# PSAppendix - Selected Primary Studies

### v1

@@ 4.2.2. Data collection
Some studies have more exploratory data collection steps. For example, data collection is combined with data discovery and data selection, i.e., collecting data for a drug discovery task requires performing experiments to produce data and selection of data by domain experts to be used in later phases [37]. On the other hand, there are cases that follow a prior data management plan prepared by data engineers to collect data to be fed into the ML pipeline [117]. For security-critical systems, data collection process is applied with anonymization and encryption of the collected information to ensure data privacy [31].
4.2.3. Data preparation
This step aims to prepare data for the ML model building phase, by processing the data including cleaning the raw data, data augmentation, labeling data points, data transformation, and feature engineering according to the requirements of the ML model.
Data cleaning refers to detecting and fixing errors in raw data, e.g., duplication removal [24], removing invalid and irrelevant data points [27], [111], imputing missing values in data [37], [117], [111]. This step is also referred to as data filtering in the literature [27].
Data augmentation is an activity that helps to increase the data sample size to overcome problems that might occur due to limited data availability. Applying oversampling and using simulation are two techniques mentioned [5].
Data labeling is the task of marking data points according to the observations in the problem domain. Data engineers, data scientists, and domain experts take part in this task [117]. While data labeling is a manual task in many cases, there are tools that support automated labeling [111].
Data transforming means adjusting the format of the data in accordance with the needs of the ML system to be built, e.g., removing some characters to make the data content available for use in other file systems [27]. This step also includes converting data into another representation format, i.e., applying distributional transformation techniques on numerical data [31], and applying data discretization for converting the data into smaller units [31].
4.2.4. Data analysis
Data analysis aims to explore the data statistics, e.g., the mean and standard deviation in data, quantiles of data [30], [62] to deeply comprehend the data before the model development step. It helps to understand and assess which features [112] and data slices [30] are best for the ML model requirements. Data analysis also offers data visualization using dashboards to the team to make their evaluations of data easier [25], [42], [114], [109]. Moreover, data analysis helps data validation, which aims to assess the health of data, i.e., data might involve anomalies or unexpected properties that might affect the model quality [62].
Feature engineering focuses on processing features in data to create successfully performing ML models, i.e., improving the ML model’s prediction performance and outcome accuracy [64], [134], [112]. In the literature, feature engineering is mentioned with respect to the following sub-steps: feature transformation, feature selection, feature extraction, feature construction, and feature storage [162], [114], [115], [102]. Our analysis shows that feature engineering is mostly seen as the responsibility of data engineers and scientists. Also, as with many data-related MLOps activities, feature engineering uses the expertise of other roles such as security [105] and product experts [31].
Another mentioned activity in the literature that falls under data discovery through data analysis is data slicing [137]. Data slicing is the process of finding the subset of data that performs according to the use case. It overlaps with the model building and evaluation steps since it explores the interest in data through evaluating the ML model’s prediction performance on different data slices.
4.2.5. Model development
This activity is the traditional step in ML life-cycle that aims to construct the ML model(s). Model building includes model design, training, evaluation, selection, validation, and model re-training activities. The model building partially overlaps with data analysis since the building of an ML model often involves exploring multiple training approaches, i.e., ML algorithm and hyperparameter configurations [114].
Model design refers to designing the appropriate architecture or selecting the appropriate tools for ML model construction to ensure successful model training and prediction performance. Examples are designing the architecture of neural network models [39], selecting the right algorithm for the problem domain [24], and designing the best performing ensemble [31].
Model training is building an ML model that will often be used for prediction. Traditionally in the ML life-cycle, the model training step takes the prepared dataset and by utilizing a learning algorithm it creates a prediction model learned from the prepared data [111], [103], [120], [125].
Model evaluation is assessing the predictive performance of the ML model. This step is traditionally done by testing the ML model with a held-out test set, which is a separate set of data points that are not used to train the model [29; 139; 141; 114; 102]. The model’s performance on the test set is evaluated using appropriate metrics for the use case [120].
Model optimization refers to improving a model to make it more successful in terms of performance. One of the most common approaches used in the literature is hyperparameter tuning [24], [102].
Model selection step deals with deciding the most appropriate model for the problem domain among various models built for the same ML system [1], [7], [152].
Model validation step is the evaluation of the model according to the business goals [70], [110], quality standards [62], [114], and regulatory requirements [110], [114]. Deployment of a model is decided based on the model validation results. Model validation often requires the supervision of a manager [34], and the collaboration of data scientists, business stakeholders, and ML engineers [114]. The common method for validating the quality of an ML model is comparing its predictive performance against a baseline model or a fixed threshold value [62], [157], [102].
@@ 4.2.6. Deployment
This is the process of deploying trained ML models into production environments, where the model predictions are accessible to users. Depending on the use case, deployment is done by integrating an ML model into a software application or serving the model as a service.
As reported in [117], ”the deployment stage is a pivotal moment in the MLOps process”, mainly due to the differences between the phases prior to the deployment and after the deployment. The majority of the tasks prior to the deployment phase correspond to the traditional ML tasks, i.e., data collecting, data preparing, and model building, that are under the responsibility of data and ML-related roles (i.e., data engineers, data scientists, and ML engineers). However, deployment related tasks differentiate from the routine ML activities [143], mostly include model packaging [138] and engineering software infrastructures, e.g., deployment in cloud [102]. Deployment is the phase in which engineers from different backgrounds, i.e., ML engineers and software engineers, are involved in the process. In prior phases, engineers and data scientists mostly work experiments in a research environment. During the deployment phase, it is software engineers and MLOps engineers working on the smooth integration of the trained ML model into a running production environment [117], [110], [108].
The deployment phase addresses various aspects of ML operations, i.e., model packaging to share the trained models with the community [24], packaging and creating an image of the model that also includes its dependent libraries [70], [141], [144], [139], [138].
Deployment methodologies and complexity of the deployment vary in the literature depending on the target ML system [24], [127]. A common way of making the predictions of an ML model available to users is providing an application that enables users to query the ML model and serve its predictions [29]. Many studies prefer the model as a service approach which allows serving the model through REST API end point [29], [37], [141].
@@ 4.2.7. Monitoring
Monitoring is one of the key activities of MLOps that is applied by many studies and supported by many tools [26], [32], [152],[114], [109], [148]. It helps to improve the system’s health through receiving feedback which often leads to repetition of the necessary ML tasks.
Monitoring activity aims to ensure the service quality and operational stability of the ML system in production by continuously observing the system resources, dependencies, and performance to capture the issues in a timely manner [145], [117]. The task results in a notification of the situation of the system [109] and often leads to an invocation of earlier tasks in the pipeline [102], [114]. Besides the detection of erroneous cases, monitoring is also used to gain perspective on how pipeline tasks are running and how accurate the model’s performance is [91]. Monitoring also helps to increase users’ and business stakeholders’ trust by enabling system observability [26], [146], [100].
The most commonly monitored subjects in ML systems are concept drift, quality of data and model, performance of the model, usage of computational resources such as disk usage, and non-functional requirements of the system such as security [24], [32], [114], [111], [148].
Monitoring is under the responsibility of MLOps engineers [137], [111]. However, in many cases, other roles are involved in monitoring, i.e., product performance experts, site reliability engineers, and solution engineers consume the output of the monitoring to examine the issues (PS, 31).
Systems and organizations that are at a more mature stage e.g., highly regulated industries like financial services [99], are also concerned with the explainability of ML models as part of monitoring operations [99], [25]. Explainability helps a better understanding of the model’s predictions, i.e., visualization of the models’ decisions [17], and providing a rule-based root-cause analysis [23].
@@ 4.2.8. Pipeline management
An MLOps pipeline is a flow of traditional ML life-cycle tasks and ML operations. These pipelines can easily become complex since there is a variety of activities that connect with each other through the MLOps life-cycle. Management of pipelines aims to execute MLOps tasks efficiently and reliably by addressing pipeline design, orchestration, and optimization.
Pipeline automation aims to reduce human intervention to a minimum degree by enabling the execution of MLOps tasks automatically and easing the handling of repeating complex ML tasks and operations [42]. Many studies work on the automation of MLOps tasks through a pipeline designed according to their use case. For example, one study focuses on the automated pipeline of data downloading, data processing, model training, and testing steps [3], and another focuses on the automated pipeline of the model configuration, serving, and monitoring [44].
Automation requires pipeline triggers to enable the automated repetition of the pipeline tasks. At the most basic level, triggers are set to execute a task in the pipeline when the prior one is completed [42]. But pipelines can also be triggered based on multiple factors depending on the use case [102]. Tasks in the pipeline can be triggered periodically according to a scheduled period [25], [157], [152], deployment can be triggered when a newly trained model is available [148], the training task can be triggered when new data is available or the model performance is reduced [141], [69], training can be triggered when the distribution of data significantly changes (concept drift) [102].
Scaling pipelines has importance in today’s infrastructures. More specifically, ML-enabled systems work on edge devices, and many ML systems and tools are served to different customers. So, the same pipeline structure may not fit all. Scaling and adjusting the complexity of pipelines is important. Keeping data pipelines scalable to manage the evolving customer requirements is proposed [42], Airflow supports the scalability of pipelines to handle complex and dependent structures [153].
@@ 4.2.9. Artifact management
Versioning the code, data, models, experiments, metadata, and documents is an important practice in ML projects as it is in traditional software engineering projects. The aims of this activity involve keeping a record of history and versions which allow easier collaboration among team members [33], [36], debugging the errors and backing up the system [157], [102], artifact reuse [32], creating an audit trail for quality and risk assessments [24], [134], providing transparency [41] in the project.
Specifically, studies focus on versioning data science experiments [122], [108], [100], codes [24], [36], ML models [95], [65], training metadata [134], parameter weights used for model training [157], tracking software versions running with the ML model [41], [42], logging ML pipeline executions [157], [102], logging monitoring events [33], and logging container data [90]. Further, data [24] and model [190] provenance are stored which are the records that keep the origin and lineage of data and model in ML enabled systems [66].
@@ 4.2.10. Quality assurance
The quality of ML models and systems through their lifecycle is important. Quality assurance covers data quality and model quality, as well as covering the overall system’s quality. Data quality addresses accessibility, completeness, consistency in data format and structure, accuracy (e.g., realistic values), data heterogeneity, fairness of collected data [24], [152]. Model quality mostly addresses the assessment of the predictive performance of the models through the evaluation of performance metrics and test set [21], [107]. Model quality assessment often results in model optimization. Model quality also refers to bias and fairness [100], [152].
4.2.11. Collaboration
As MLOps become popular in recent years, organizations needed new collaboration practices for smoother communication across people from different backgrounds (i.e., ML engineers and software engineers) to manage the MLOps projects smoothly [151]. One of the important aspects regarding the collaboration theme is identifying collaboration points for MLOps teams to organize the communication between roles from different backgrounds [151]. These collaboration points ease communication across different people.
4.3. RQ1.3 - What are the Practitioner Roles and Responsibilities in MLOps
4.3.1. Data Engineer
Mainly, data engineers are responsible for fundamental data activities and enabling data operations in a high-quality, secure, and scalable way [1], [24], [42], [85].
In a typical MLOps project, data engineers’ main set of tasks are identifying suitable data sources, checking the legality of data, collecting data from various sources, cleaning them, and preparing the labeled data [5], [51], [93]. Here, data engineering work is characterized as ”extract-transfer-load” process [112], [106]. It also involves transferring data into convenient formats to make it consumable by data scientists and ML engineers [112], [106]. The essential part is gathering and preparing the data for further steps which are data analytics, ML engineering, and operations. These steps are often executed in collaboration with domain and business experts, data scientists, and ML engineers.
Some organizations have specialized roles within their teams whose responsibilities overlap with a typical data engineer’s responsibilities. For instance, data curator provides data and implements custom data acquisition adapters [19] while data expert applies specific noise reduction techniques to prepare data for further steps [5].
The responsibilities of data engineers are not limited to basic data-collection and preparation activities, they are also involved into data operations, i.e., setting up data storage infrastructures and building automated data pipelines [1]. Ensuring that the ML system’s infrastructure is right for data storage and other data operations is managed by data engineers [69]. Leading by the size of data and structure of the system, data engineers build data storages and pipelines, e.g., storing data curated from distributed sources in a centralized data lake [69], enabling big data streaming platforms [125]. Designing and constructing automated data pipelines is critical for making data accessible to the MLOps team for ML operations at scale [72], [85], [106], [112], [109].
Data engineers are also responsible for security controls of data, data privacy, data quality, and governance [86], [106]. While data security, quality, and governance responsibilities are shared among many roles, i.e., data engineers, MLOps engineers, domain and business experts, and security experts, there are two specific roles defined for mainly these purposes; data steward who is responsible for data quality management [24], [117], and data governance officer who ensures data governance and privacy [114].
4.3.2. Data Scientist
Data scientists are mainly responsible for exploring and analyzing the data made available by data engineers and then training, tuning, and evaluating a model to derive insights from the data [122], [124], [187], [161], [104].
In a typical MLOps project, the initial responsibility of a data scientist is to understand the business problem and its domain, and then evaluate the feasibility of solving the problem through building an ML-enabled system [24], [114]. At this step, data scientists work with domain and business experts to discuss objectives and key results (OKRs) [86]. During this step, experiments conducted by data scientists play a significant role in evaluating the suitability of the collected data to address the business problem, as well as evaluating the effect of feature engineering and model design for more feasible ML solutions.
Feature engineering is often conducted by data scientists to analyze the best combination of features in data to train more feasible ML models. Later, data scientists conduct model training and evaluation tasks to generate various models to select the most appropriate, e.g., the best-performing, one among them. During this stage, dataset splitting into training, test, and validation sets, choosing the suitable ML algorithm, training the model, choosing suitable model evaluation metrics (e.g., accuracy), evaluating the model accordingly, and selecting the suitable model are the responsibilities of data scientists.
These responsibilities of data scientists necessitate collaboration with other roles, namely business experts and ML engineers. For example, during model validation, data scientists need to assess whether the trained model fits technical, business, or regulatory requirements [114].
Experiments conducted by data scientists typically iterated through multiple iterations until a defined goal (e.g., accuracy threshold for the model) is reached [24]. Writing experiment codes and ensuring the trackability of multiple versions of these codes, datasets, and models are the responsibility of data scientists.
Moreover, as data engineers are responsible for the initial part of an ML pipeline, data scientists are responsible for designing and building scalable and high-quality data analytics pipelines for experimentation and model training. A data scientist’s last responsibility in the ML pipeline is early monitoring [114], [93]. When a model is deployed, monitoring its early results is often accompanied by data scientists as well as ML and MLOps engineers, since the early results may not be as expected which often leads to returning to the feature engineering, model training, and validation steps.
4.3.3. ML Engineer
ML engineer and data scientist roles may overlap, or sometimes these roles are mentioned interchangeably [135], [141]. However, many studies distinguish definitions of both roles [114], [154]. According to those, an ML engineer’s task is primarily model optimization, deployment, and monitoring.
ML engineers’ tasks are experimenting with datasets [118; 106], experimenting with models [118], and optimizing the model and its parameters [24], [6], [118], [107], [108]. These are also the responsibilities of a typical data scientist in an MLOps project.
In MLOps projects, ML engineers’ responsibilities are not limited to these: They contribute to the monitoring and production activities [114], [107]. ML engineers collaborate with data scientists and business stakeholders to ensure the model selected for deployment confirms the necessary level of performance and business requirements. ML engineers take responsibility in the deployment and monitoring stages of the ML pipeline. ML engineers take responsibility for executing a suitable deployment for the business use case, e.g., online serving, and monitoring the deployed model’s performance in the production environment.
4.3.4. MLOps Engineer
MLOps engineers are responsible for building automated pipelines, maintaining the operational stability of the entire ML system, and collaboration across different roles [106], [117], [151].
An automated pipeline for data collection, data analysis, model development, and deployment phases is designed and practiced by data engineers, data scientists, and ML engineers respectively. However, the main responsibility for building and optimizing these automated pipelines belongs to MLOps engineers. Particularly, MLOps engineers integrate different services to pipelines (e.g., cloud, container), maintain the management of dependent packages, versioning and tracking mechanism of the project, and overall platform governance [106].
Monitoring the data, model, and application is often carried out by MLOps engineers along with ML engineers [117]. MLOps engineer steps in when human intervention is needed during monitoring, i.e., model needs re-training in the case of data shift, dependencies between data and model need to be resolved [105], [137]. All these tasks require managing the communication across different roles in an MLOps project [105].
4.3.5. Software and application developers
take crucial responsibilities in many MLOps teams, e.g., [144], [104]. They develop the final applications and services that help ML production. In many cases, they collaborate with data scientists and ML engineers in analyzing and processing data, and building models [115], [151]. Besides these, software engineers are also responsible for the deployment, packing, releasing, configuring, and monitoring of ML systems [161], [144], [154]. Further, the security of the ML system is maintained by software engineers specialized in security [105].
4.3.6. Operational engineers
focus on mainly monitoring activities [104], while they also take roles in deployment operations, i.e., packaging, releasing, and configuring of ML systems [144].
4.3.7. Technical and solution architects
are responsible for choosing the right technologies and design paradigms for building ML systems, integrating ML systems into cloud-based applications, and ensuring the building of secure ML systems [105], [155].
4.3.8. Quality, governance, risk, and compliance roles
are not typically engaged in day-to-day MLOps operations [114], as their primary tasks revolve around quality assurance, risk management, and governance. Studies define various roles for quality and governance across multiple MLOps aspects such as data, model serving, and model performance, etc. Particularly, data governance officer, data steward, and data quality board are the roles mentioned in the literature for being responsible for data governance, data quality management, and data privacy [117], [114]. Compliance teams are committed to mitigate unfairness and bias in ML systems across model training, deployment, and monitoring phases [108], [152]. Recently, risk professionals become involved in MLOps projects to manage data and model-related risks [155]. Specifically, model risk manager is responsible for ML model inventory management, model explainability, and pipeline management [111], [106], while other studies only mention risk professionals for managing any data and technology related risks [155].
4.3.9. Domain experts

### v2

@@ 3.2.2. ML lifecycle activities
Data collection is the first step in the ML pipeline, involving the acquisition of raw data that will be processed in later steps to train a ML model. Often, raw data is gathered from relevant data sources, e.g., sensors [24], GitHub code repositories [27], or cameras [34]. Some studies have more exploratory data collection steps. For example, data collection is combined with data discovery and data selection, i.e., collecting data for a drug discovery task requires performing experiments to produce data, followed by selection of data by domain experts to be used in later phases [37]. Alternatively, some cases follow a data management plan prepared by data engineers to collect data to be fed into the ML pipeline [117]. For security-critical systems, data collection process is applied with anonymization and encryption of the collected information to ensure data privacy [31].
Data preparation aims to process raw data for ML model building through cleaning the raw data, augmentation, labeling data points, transformation, and feature engineering. Data cleaning involves detecting and fixing errors in raw data, such as removing duplicates [24], filtering invalid and irrelevant data points [27], imputing missing values in data [37]. This step is also known as data filtering in the literature [27]. Data augmentation increases sample size to address limited data availability, with techniques like oversampling and simulation [5]. Data labeling marks data points according to observations in the problem domain, involves data engineers, data scientists, and domain experts [117]. While data labeling is a manual task in many cases, tools also support automated labeling [111]. Data transforming adjusts the data format according to the needs of the ML system to be built, for example, removing characters for compatibility with other file systems [27]. It also includes data representation changes, like distributional transformations on numerical data and data discretization for smaller units [31].
Data analysis includes data exploration based on data statistics, e.g., the mean and standard deviation, and quantiles[30], [62] to better understand the data before model development. It helps to understand and assess which features [112] and data slices [30] best meet ML model requirements. Data analysis also includes data visualization providing dashboards to facilitate the team’s evaluations of data [25], [42], [114], [109]. Moreover, data analysis supports data validation, which assesses the health of data by identifying anomalies or unexpected properties that could impact model quality [62]. Feature engineering processes features to enhance model performance and accuracy [64], [134]. It includes sub-steps such as feature transformation, feature selection, feature extraction, feature construction, and feature storage [162], [114], [115], [102]. Feature engineering is mostly seen as the responsibility of data engineers and scientists but also uses the expertise of other roles such as security [105] and product experts [31]. Another mentioned activity is data slicing [137], which is the process of finding the subset of data that performs according to the use case. It overlaps with model building and evaluation steps since it explores the interest in data through evaluating the ML model’s prediction performance on different data slices.
Model development is a core step in ML life-cycle, focused on constructing ML models. It includes model design, training, evaluation, selection, validation, and re-training. Model building partially overlaps with data analysis as it often involves exploring various training approaches, i.e., ML algorithm and hyperparameter configurations [114]. Model design refers to designing and selecting the appropriate architecture and tools for ML model construction to ensure successful model training and prediction performance. Examples include designing neural network architectures [39], choosing the right algorithm for the problem [24], and optimizing ensemble models [31]. Model training builds the ML model for prediction, using a learning algorithm to train it on the prepared data [111], [103]. Model evaluation assesses predictive performance of the ML model, by testing it with a held-out test set, which includes a separate set of data points not used in training [29].Prediction performance is measured using relevant metrics for the use case [120]. Model optimization improves the model’s prediction performance. Hyperparameter tuning is a common method applied [24], [102]. Model validation evaluates the model based on business goals [70], quality standards [62], and regulatory requirements [110]. The common method for validating the quality of an ML model is comparing its predictive performance against a baseline model or a fixed threshold value [62], [157], [102].
3.2.3. Application development
step addresses the implementation of software and/or service that will be interfaced with the deployed ML solution [108], [136]. Additionally, this step involves verification, validation, and quality assurance activities of the developed application such as unit testing, integration testing, refactoring, maintenance to ensure reliability, functionality, and seamless interaction with the ML model [117], [124].
@@ 3.2.4. Deployment
The deployment phase integrates trained ML models into production environments, making model predictions accessible to users. Depending on the use case, deployment is achieved by integrating the ML model into a software system [117], serving it as a service [14], or creating containerized images of ML models [141]. Deployment phase differs from routine ML activities [143], as it involves transferring the model to another team or infrastructure [110]. As Neghawi et al. [117] note, ”the deployment stage is a pivotal moment in the MLOps process,” due to the transition from an experimental (ML development) to an actualized environment (ML production). Key deployment activities include model selection for choosing the most appropriate model among various models built for the same ML system, i.e., the best performing one [180], or based on client’s requirements [142]. Deployment of a model is decided based on model validation results, which often involve supervision of a manager [34], and collaboration of data scientists, business stakeholders, and ML engineers [114]. Another key activity is model packaging for reproducibility and sharing [24], creating containerized model images with dependencies for inference [70] , and selecting deployment methodologies based on the target ML system [24], [127]. Model serving refers to making the ML model accessible for generating new predictions. A common approach is embedding the model within an application for user queries [29]. Many studies favor the Model-as-a-Service paradigm, where models are exposed through REST API endpoints for scalability and ease of access [29], [141]. Deployment requires collaboration from ML and MLOps engineers, as well as software engineers, to ensure smooth integration into a production environment.
@@ 3.2.5. Monitoring
Monitoring is a key MLOps activity applied by many studies [26], [32] and supported by many tools [152],[114], [109], [148]. It helps improve system health by providing feedback, often triggering the repetition of relevant MLOps activities. Ensuring operational stability and quality of an ML system in production is done by continuously observing data and model changes, model performance, system resources, application and its dependencies to capture and report the issues in a timely manner [145], [117]. Input data changes, data and model biases, and concept drifts can be detected setting pre-defined thresholds and rules on data [111], or monitoring the model performance [148]. When a change, bias, or drift is captured through monitoring, it often results in adjusting the dataset for model training and/or model optimization [190]. Captured model performance degrades invokes model re-training [114] or relevant ML processes [141]. Other common monitored subjects are resource usage, e.g., memory usage and number of API calls [24], non-functional requirements like security [32]. Monitoring changes around system dependencies helps providing notifications regarding specific modifications [109]. Information collected during monitoring is stored and used to produce notifications and reports of important situations which often leads to an invocation of earlier tasks in the pipeline [152]. Dashboards [114] and reports [102] are common methods for presenting this information. Moreover, monitoring enhances users’ and business stakeholders’ trust by enabling system observability [146], e.g., providing insights on how pipeline tasks are executed [91]. While MLOps engineers are primarily responsible for monitoring [137], [111], other roles, such as product performance experts, site reliability engineers, and solution engineers, also utilize monitoring data to analyze issues (PS, 31). In more mature systems and organizations, such as those in highly regulated industries like financial services [99], monitoring also includes explainability features to enhance understanding of model predictions, including decision visualizations [17] and rule-based root-cause analysis [23].
@@ 3.2.6. Pipeline management
An MLOps pipeline consists of interconnected ML life-cycle tasks and operations, which can become easily complex due to their dependencies. Pipeline management addresses pipeline design, orchestration, and optimization to execute MLOps tasks efficiently and reliably. Pipeline automation minimizes human intervention by executing MLOps tasks automatically, simplifying repetition of complex ML operations [42]. Studies have explored automation for different use cases, such as pipelines for data downloading, processing, model training, and testing [3], or for model configuration, serving, and monitoring [44]. Automation relies on pipeline triggers to enable task execution. Triggers can be sequential, i.e., executing a task upon the previous task’s completion [42], periodical e.g., executing according to a pre-scheduled period [25], event-driven, e.g., deployment can be triggered when a new ML model is available [148], e.g., training can be triggered when new data is available or the model performance is reduced [141][141], or a concept drift is detected [102]. Pipeline automation is often tailored to specific use cases. For instance, one study automates data downloading, processing, model training, and testing [3], while another focuses on automating model configuration, serving, and monitoring [44]. Moreover, scaling pipelines is essential in today’s infrastructures such as diverse environments including edge devices and multi-tenant platforms. For example, data pipelines must adapt to manage the evolving customer requirements [42], tools like Airflow support scalable, dependency-aware pipelines [153].
@@ 3.2.7. Artifact management
Versioning artifacts is crucial in MLOps projects just as in traditional software engineering. Common MLOps artifacts are code, data, feature, models, experiments, metadata, and project documents. Managing artifacts involves keeping a record of history and different versions of artifacts. It facilitates collaboration among team members [33], enables debugging and system backups [157], supports artifact reuse [32], ensures auditability for quality and risk assessments [24], [134], and enhances project transparency [41]. Studies focus on versioning various MLOps components, including data science experiments [122], codes [24], ML models [65], training metadata such as dataset versions [134], parameter weights used for model training [157], tracking software versions running with the ML model [41], logs of pipeline executions [157], logs of monitoring events [33], and logs of container data [90]. Further, data [24] and model [190] provenance are stored which are the records that keep the origin and lineage of data and model in ML enabled systems [66].
@@ 3.2.8. Quality assurance
Quality assurance covers ensuring data quality, model quality, and ML application’s quality. Quality is managed across various MLOps phases, but data quality is closely associated with data cleaning, analyzing data, model training, model selection, and monitoring phases [24]. Model quality is managed through data collection, model training and evaluation, and monitoring phases [111]. Moreover, ML application’s quality is ensured through traditional software testing strategies [100]. Data quality involves assessing the scientific quality of data with domain experts, e.g., whether data includes realistic values, correcting data anomalies, checking data completeness, e.g., addressing problem-related missing values, checking consistency in data format and structure, data heterogeneity, and fairness of collected data [62], [24], [152]. Model quality involves assessing the model’s predictive performance through performance metrics and test sets [21], validating model bias and fairness [100], maintaining ML portfolio health [90], monitoring model to detect overfitting issues [59], dimensionality reduction [24]. Application quality involves unit testing, acceptance testing [89], maintaining containerized images and identifying vulnerabilities [90].
3.2.9. Team management
Various people from different backgrounds, such as ML engineers and software engineers, need to work together in MLOps settings [151]. Facilitating communication among these individuals is crucial to maintaining collaboration and ensuring smooth operation of ML system to achieve business goals [105]. Activities associated with team management include clear identification of requirements, planning the project, ensuring transparency in project documentation and model evaluation, involving individuals with relevant expertise in meetings and negotiations, and investing in education of team members [151].
3.3. RQ1.3 - What are the Practitioner Roles and Responsibilities in MLOps?